OpenStudy (anonymous):

Each side of a square is increasing at a rate of 6cm/s. At what rate is the area of the square increasing when the area is 16cm?

OpenStudy (anonymous):

\[A = l^{2}] dA/dt = 2*l (dl/dt) dA/dt = 2*4*(6) = 48 cm^2/s